# D-LiFT: Improving LLM-based Decompiler Backend via Code Quality-driven Fine-tuning

### v1

@@ 3 Design Challenges
Challenge 1: LLM Improvement Method Selection. Improving the code generation capability of LLMs [32, 54, 55, 45, 15] is not new. However, many of them are challenging to adapt for decompiled code improvement. Unlike typical code improvement tasks, decompiled-code enhancement faces a unique challenge: the existence of multiple valid ground truths. Specifically, a single binary may be compiled from multiple semantically equivalent source programs, each of which should be considered a valid ground truth. For instance, as illustrated in Figure 1, two different source files, each around 300 characters long and with an edit distance of 74, produce the same binary output using GCC [1] with -O2 optimization. In practice, however, researchers typically have access to only one of these source variants as the ground truth. Moreover, many fine-tuning methods, such as supervised fine-tuning (SFT) and the actor-critic paradigm in RL, accept only a single reference to compute the training loss, overlooking the possible existence of a full set of correct alternatives. This constraint may degrade a model’s effectiveness, as exemplified by LLM4Decompile’s use of SFT on the source code during training.
Challenge 2: Assessment of decompiled code quality. The assessment of decompiled code quality should reflect the principle of preserving decompiled code accuracy while improving readability. To the best of our knowledge, there is no existing code quality assessment function that addresses these specific requirements.
@@
For accuracy, as introduced in Section 2.2, we mainly focus on two aspects, syntactic correctness and semantic correctness. Since the assessment function must be deterministic, methods such as fuzzing  [68] and its derivatives, like MSSC [11], are excluded. Additionally, to ensure applicability without the need for source code, we exclude approaches that rely on source-dependent checks, such as unit tests or Alive2 [39]. That leaves D-helix [72], which uses an iterative re-compiler to examine the syntax errors and symbolic execution to check the semantics against the original binary. However, integrating D-helix directly into our framework introduces its own challenge. Notably, D-helix still relies on certain decompiled code output artifacts, e.g., external function calls, to analyze. Since LLM-generated code frequently omits instructions, these artifacts may be missing or unreliable, undermining D-helix’s effectiveness.
@@
For readability, existing metrics [10, 42, 47, 51] each have limitations, e.g., ignoring decompiler-specific artifacts, when applied directly to decompiled code. To the best of our knowledge, R2I [17] is the only metric tailored specifically for decompiled output. However, since it’s a relative measure, whose features are derived from multiple decompilers, it cannot assign an absolute score to standalone LLM-generated code. For example, the feature, “number of unnecessary goto labels”, is calculated by contrasting angr [61] or RetDec [5] outputs against those from Ghidra [3] and Hex-Rays [2], which cannot be computed when only one decompiler’s output is given.
@@
In summary, our principle, preserving accuracy while improving readability, cannot be satisfied by any single metric currently available.
4 Design
@@
Figure 2: This figure shows how D-LiFT fine-tunes the policy model (i.e., the baseline model). Specifically, it applies GRPO to overcome the inherent complexity of decompilation tasks, where single decompiled code segments can correspond to multiple semantically equivalent source representations.
@@
D-LiFT is an automatic decompilation pipeline that consists of a decompiler front-end and an LLM back-end capable of improving the quality of decompiled code. Figure 2 illustrates the workflow by which D-LiFT fine-tunes an LLM to produce improved decompiled code output. Three inputs are needed: a policy model (i.e., LLM waiting for training), a set of training binaries, and a decompiler. To improve the policy model’s capabilities of generating better decompiled code, D-LiFT employs reinforcement learning (Section 4.1). To guide this training, D-LiFT introduces a multi-aspect assessment/reward metric, D-Score, that evaluates the LLM’s output code on both accuracy and readability (Section 4.2). Specifically, for accuracy, D-Score adopts D-helix’s stepwise approach: first verifying syntax (Section 4.2.1), then validating semantics against the original binary (Section 4.2.2). Only if a candidate passes these accuracy checks does D-Score move on to assess readability (Section 4.2.3). Once the training is complete, D-LiFT outputs a fine-tuned LLM capable of generating decompiled code with improved accuracy and readability.
4.1 LLM Fine-Tuning
@@
The challenge described in Figure 1 highlights a critical limitation in current approaches: decompiled code can map to multiple semantically equivalent source representations, each representing a valid ground truth. This multiplicity poses significant training difficulties for supervised fine-tuning (SFT) approaches, such as LLM4decompile, which are designed to work with only one correct answer. Although reinforcement learning generates multiple candidates during training, theoretically providing access to various valid solutions, conventional RL policies maintain the single ground truth constraint when computing the loss. Fortunately, Group Relative Policy Optimization (GRPO) [53] overcomes this constraint by generating the loss from the normalized rewards across multiple candidate outputs. ,
Figure 2 illustrates how D-LiFT integrates GRPO into our workflow. WAH1g1 Given a training binary, denoted as $bin$, D-LiFT invokes the heuristic decompiler to produce a original decompiled output, denoted as $o_{g}$. WAH1g2 Using the $o_{g}$ as input, the policy model generates a set of candidate refinements,$\{o_{1},o_{2},...\}$. WAH1g3 For each candidate, $o_{i}$, D-LiFT computes a reward, $r_{i}$, using D-Score (see Section 4.2 for details). WAH1g4 Next, GRPO normalizes these rewards, $\mathbf{r}$ = $\{r_{1},r_{2},...\}$ within each candidate group according to its standard formulation:
@@
WAH1g5
Finally, GRPO uses these normalized rewards, $\hat{A}_{i,t}$, to compute the loss for fine-tuning the model as follows:
@@
Given that the decompilation task accepts multiple valid ground truths, D-LiFT eliminates the KL-divergence weight by setting $\beta$ to zero, thereby removing dependence on a single reference solution. In practice, setting $\beta$ to zero not only reduces memory and computational overhead but also improves training effectiveness, in line with recent findings [24, 38]. The additional variables referenced in Equation 2 serve to modify the policy model parameters, with detailed definitions available in the original GRPO research publication [53].
4.2 Reward Function Design
Figure 3 shows the overall procedure of D-Score. As a reward function, D-Score takes three inputs: the original decompiled code, $o_{g}$, the original binary, $bin$, and a candidate refinement from the policy model, $o_{i}$, and outputs a comprehensive score $r_{i}$ reflecting both accuracy and readability of $o_{i}$. Formally, $r_{i}$ is defined as follows:
@@ 4.2 Reward Function Design
[…]
Within the above equation, $X_{pen}$ represents the penalty score because of error $X$, and $R_{Y}$ represents the reward return from each specific check in field $Y$. Note that, following our principle, we also define:
$syn_{pen}<ret_{pen}<call_{pen}<min(R_{read})$ (4)
To generate $r_{i}$, specifically, WAH1g1 D-Score checks for syntax errors (See Section 4.2.1). If any is found, it returns a syntax error score $syn_{pen}$. WAH1g2 Otherwise, it verifies semantic equivalence between $o_{i}$ and $bin$. If $o_{i}$ fails the semantic check, D-Score returns a score, $R_{sem}$, based on the symbolic matching with the $bin$ (See Section 4.2.2). WAH1g3 Only when $o_{i}$ passes both syntax and semantic checks, D-Score defers to the readability component and returns a readability score,$R_{read}$ (See Section 4.2.3).
Figure 3: D-Score Framework, following our principle of preserving accuracy while improving readability.
4.2.1 Syntax Metric
Since D-LiFT uses D-helix’s framework to examine accuracy, D-Score adopts D-helix’s Recompiler. As an iterative recompilation tool, Recompiler automatically initializes undefined variables, injects required system libraries, and translates special pseudo-instructions (e.g., CONCAT) into function calls. For each candidate code $o_{i}$ from the policy model, D-Score submits it to Recompiler and obtains a syntax score, $R_{syn}$, based on Recompiler’s feedback. Specifically, D-Score assign $R_{syn}(o_{i})$ to $syn_{pen}$ when the $o_{i}$ cannot be recompiled and 00 otherwise. Formally:
$R_{syn}(o_{i})$ =$=\left\{\begin{array}[]{l}syn_{pen},\text{ if }o_{i}\text{ cannotbe compiled. }\\ +0,\text{ otherwise }\\ \end{array}\right.$ (7)
@@ 4.2.1 Syntax Metric
Note that $syn_{pen}$ contributes directly to D-Score (as shown in Equation 3), ensuring a clear penalty for any syntax error.
4.2.2 Semantics Metric
@@
Figure 4: The workflow of semantic check in D-Score. The left side shows how the return value is checked, and the right side shows how the external function call is examined.
@@
Once the $o_{i}$ passes the syntax check, D-Score runs symbolic execution to conduct the semantic check between the original binary and $o_{i}$’s IR. In addition to the behavior of return values, we also consider the behavior of external function calls in our semantic equivalence check.
@@
For the return values check, we mostly follow D-helix’s approach, which employs symbolic execution and SMT solver. Specifically, as shown on the left side of Figure 4, given the decompiled code, D-Score symbolizes the input arguments and runs the symbolic execution to generate the symbolic model, Symbolic-Model-Ret, a mathematical model that encapsulates function behavior by mapping formula inputs to function arguments and formula outputs to function return values. After that, D-Score runs the SMT solver on SMT-RET to compare the symbolic models between the original binary and the decompiled code.
@@
When checking the behavior of the external function call, however, directly applying D-helix to test LLM-generated code (e.g., $o_{i}$) can yield many false negatives. This limitation stems from D-helix’s fundamental dependence on decompiled code to establish ground truth for verifying external function calls. Specifically, it models external function calls by approximating return values through the sum of the least significant bytes of those arguments. However, this external function call modeling process is only initialized when an external function call is detected within the decompiled code. Since instructions, including function calls, are often omitted in the LLM-generated code, D-helix may not properly initialize its function call modeling procedures, leading to possible false negatives.
@@
To model the external function call in LLM-generated code in a better way, D-Score counts how many external functions are invoked, as shown on the right side of Figure 4. Specifically, to identify valid external calls in each decompiled function, we first extract and save the names of all invoked external functions in the original decompiled output, $o_{g}$, to a file as our ground truth. After that, during the symbolic execution of the LLM-generated code, we build a symbolic model called Symbolic-Model-Call to model the external function call behavior. Unlike Symbolic-Model-Ret tracking the return value, this model uses the ground-truth file to track the number of matching function calls in the LLM-generated code for each execution path. Finally, we run SMT solver on SMT-CALL to compare the function call-count symbolic models between the original binary and the decompiled code. Note that we ignore the arguments in our semantic checks because compilation removes this information, and existing recovery techniques  [34] cannot reliably reconstruct it.
Formally, we define our semantics metric as follows:
@@
$R_{sem}(o_{i})=\left\{\begin{array}[]{l}ret_{pen},\text{ if }(check_{ret}(o_{i})==false)\\ call_{pen},\text{ if }(check_{ret}(o_{i})==true\\ \quad\quad\quad\quad\text{\&\& }check_{call}(o_{i})==false)\\ +0,\text{ otherwise }\end{array}\right.$ (8)
@@
, where $check_{ret}(o_{i})$ is Boolean result of running SMT on SMT-RET, and $check_{call}(o_{i})$ is result of running SMT on SMT-CALL.
4.2.3 Readability Metric

### v2

@@ III Design Challenges
III-A Framework Design
As our framework is built around LLMs, we divide the challenges into two parts: training and inference.
III-A1 Training
@@
Improving the code quality of LLMs through training is not new [32, 55, 56, 46, 15]. However, many of these methods are challenging to adapt for decompiled code improvement. Unlike general code improvement tasks, enhancing decompiled code faces a unique challenge: the existence of multiple valid ground truths. Specifically, a single binary may be compiled from multiple semantically equivalent source programs, each of which should be considered a valid ground truth. For instance, as illustrated in Figure˜2, two different source files, each around 300 characters long and with an edit distance of 74, produce the same binary output using GCC [1] with -O2 optimization. In practice, however, researchers typically have access to only one of these source variants as the ground truth. Moreover, as mentioned in Section˜II-C, many fine-tuning methods, such as supervised fine-tuning (SFT) and the actor-critic paradigm in RL, accept only a single reference to compute the training loss, overlooking the possible existence of a full set of correct alternatives. This constraint may degrade a model’s effectiveness, as exemplified by LLM4Decompile’s use of SFT on one source code only.
III-A2 Inference
In scenarios where we have multiple enhanced code outputs, from the decompiler itself or various LLMs, an objective and quantitative metric is essential to rank their quality. Unfortunately, the only existing framework, DeGPT, offers no automated readability metric, relying instead on human participants’ assessments.
III-B Assessment of Decompiled Code Quality
As noted above, the existence of multiple valid ground truths (during training) and enhanced candidates (at inference) necessitates a quantitative code-quality metric to evaluate each candidate. To the best of our knowledge, no existing metric meets the dual requirement of maintaining syntactic and semantic accuracy while enhancing readability.
@@ III-B Assessment of Decompiled Code Quality
For decompiler accuracy, as introduced in Section˜II-D, two aspects of accuracy, syntactic and semantic, need to be evaluated. Since the decompiled code assessment function must be deterministic, methods such as fuzzing [70] and its derivatives, like MSSC [11] from DeGPT, are excluded. Additionally, to ensure the semantics are compared directly with the binary instead of the original source code, we exclude approaches that rely on source-dependent checks, such as unit tests or Alive2 [39]. That leaves D-helix [75], which uses an iterative re-compiler to reveal syntax errors and symbolic execution to check the code’s semantics against the original binary. However, integrating D-helix directly into our framework introduces its own challenge. Notably, D-helix still relies on certain decompiled code output artifacts, e.g., external function calls, to analyze. Since LLM-generated code frequently omits instructions, these artifacts may be missing or unreliable, undermining D-helix’s effectiveness.
@@
For readability, most existing metrics [10, 42, 48, 52] ignore decompiler-specific artifacts (e.g., the number of references/dereferences). To the best of our knowledge, R2I [17] is the only metric tailored specifically for decompiled code. However, as described in Section˜II-B, since it’s a relative measure, whose features are derived to compare between multiple decompilers, it cannot assign an absolute score to a single piece of LLM-generated code. For instance, the feature, ”number of unnecessary goto labels”, is calculated by contrasting angr [62] or RetDec [5] outputs against those from Ghidra [3] and Hex-Rays [2], which cannot be computed when only one piece of decompiled code is given.
@@
In summary, our principle, preserving accuracy while improving readability, cannot be satisfied by any single code quality metric currently available.
IV D-LiFT Design
@@
Figure 3: This figure shows how D-LiFT fine-tunes the baseline model and selects the output. It applies GRPO to overcome "single decompiled code segments can correspond to multiple semantically equivalent source representations" in training and uses D-Score to select the best quality code among inference.
@@
D-LiFT is an automatic decompilation pipeline that consists of a decompiler front-end and an LLM back-end capable of improving the quality of decompiled code. The workflow of D-LiFT, illustrated in Figure˜3, consists of a training phase (Section˜IV-A1) and an inference phase (Section˜IV-A2). Specifically, in the training phase, D-LiFT takes three inputs: a baseline model (i.e., the LLM to train), a set of training binaries, and a decompiler. It employs reinforcement learning to fine-tune the LLM, ultimately producing a model with improved decompilation capabilities. Then in the inference phase, D-LiFT first uses the decompiler to generate the initial decompiled code for an input binary. It then applies both the baseline and the fine-tuned LLMs to refine the code and selects the best-quality result among all outputs. To guide the training and inference phase , D-LiFT introduces a multi-aspect code quality assessment metric, D-Score, that evaluates a piece of candidate code with respect to both accuracy and readability (Section˜IV-B). Specifically, for accuracy, D-Score adopts the stepwise approach: first verifying syntax (Section˜IV-B1), then validating semantics against the original binary (Section˜IV-B2). Only if a candidate passes these accuracy checks will D-Score move on to assess readability (Section˜IV-B3).
IV-A System Workflow
D-LiFT works by first fine-tuning the LLM via GRPO to train the LLM with accuracy-aware rewards. Then at runtime, it will select the best-quality decompiled code for each input binary, among the outputs of the decompiler, the baseline LLM, and the fine-tuned LLM, all based on the outputs’ D-Score.
IV-A1 Training Phase
@@
The challenge described in Section˜III-A highlights a critical limitation in current approaches: decompiled code can map to multiple semantically equivalent source representations, each representing a valid ground truth. This multiplicity poses significant training difficulties for supervised fine-tuning (SFT) approaches, such as LLM4decompile, which are designed to work with only one correct answer. Although reinforcement learning generates multiple candidates during training, theoretically providing access to various valid solutions, conventional RL policies maintain the single ground truth constraint when computing the loss. Fortunately, Group Relative Policy Optimization (GRPO) [54] overcomes this constraint by generating the loss from the normalized rewards across multiple candidate outputs.
[…]
@@
and uses these normalized rewards to compute the loss for fine-tuning the model as follows:
@@
Given that the decompilation task accepts multiple valid ground truths, D-LiFT eliminates the weight related to the reference by setting $\beta$ to zero, thereby removing dependence on a single reference solution. In practice, setting $\beta$ to zero not only reduces memory and computational overhead but also improves training effectiveness, in line with recent findings [24, 38]. The additional variables referenced in Equation˜2 serve to modify the baseline model parameters, with detailed definitions available in the original GRPO research publication [54]. WAH1g5 Finally, once the training iterations are complete, D-LiFT outputs the fully fine-tuned LLM.
IV-A2 Inference Phase
As described in Section˜II-D, LLMs often introduce errors when improving code, creating a need for a quantitative metric-guided selection system. To address this, we designed the selection system as shown in Figure˜3. This system works in two steps: WAH1g1 It first takes a user’s binary file and runs a decompiler to get the original code. It then feeds that code to both the baseline LLM and our fine-tuned LLM to generate two improved versions. WAH1g2 By leveraging D-Score, the same scoring tool applied during training, it evaluates all three outputs (the original, the baseline’s, and the fine-tuned’s), selecting the highest-scoring one as the final result for the user.
IV-B D-Score Design
Figure˜4 shows the overall procedure of D-Score. As a code quality assessment function, D-Score takes three inputs: the original decompiled code from the decompiler, $o_{g}$, the original binary, binbin, and a candidate code waiting for evaluation, $o_{i}$, and outputs an absolute score $r_{i}$ reflecting both the accuracy and readability of $o_{i}$. Formally, $r_{i}$ is defined as follows:
$r_{i}$ $\mathbb$$=\mathbb{1}(\text{ cond${}_{syn}$ })\cdot\left(R_{syn}(o_{i})\right)$ (3)
@@ IV-B D-Score Design
$\mathbb$$\quad+\mathbb{1}(\neg\text{ cond${}_{sem}$ })\cdot R_{read}(o_{i},o_{g}))),$
where  $\text{\quad cond${}_{syn}$ }=\left({R_{syn}(o_{i})\ !=pass}\right),$
 $\text{\quad cond${}_{sem}$ }=\left({R_{sem}(o_{i},bin)\ !=pass})\right$.
Within the above equation, $R_{Y}$ represents the reward return from each specific check in field $Y$ and $\mathbb$$\mathbb{1}(\text{ cond${}_{x}$})$ is an "if-then-else" statement that returns 1 if the condx is True and returns 0 otherwise. Meanwhile, following our principle, inaccuracies are penalized more heavily than readability issues:
$max(R_{syn})<min(R_{sem})\;\text{and}\;max(R_{sem})<min(R_{read})$ (4)
To generate $r_{i}$, specifically, WAH1g1 D-Score checks for syntax errors (See Section˜IV-B1). If not passed, it returns a syntax penalty score. WAH1g2 Otherwise, it verifies semantic equivalence between $o_{i}$ and the original binary, binbin. If $o_{i}$ fails the semantic check, D-Score returns a semantic penalty score, based on the symbolic matching with the binbin (See Section˜IV-B2). WAH1g3 Only when $o_{i}$ passes both syntax and semantic checks, D-Score defers to the readability component, compares $o_{i}$ to the original decompiled code, $o_{g}$, and returns a readability score (See Section˜IV-B3). This overall design ensures D-Score only awards readability once the accuracy has been verified, thereby adhering to our core principle.
Figure 4: D-Score, following our principle of preserving accuracy while improving readability. It performs three sequential checks on each candidate code: syntax/semantic verification, and readability assessment. If any of the accuracy checks fail, D-Score halts immediately and returns a penalty score.
IV-B1 Syntax Metric
As prior work has shown [64, 75], decompiled output often fails to compile directly, even at the function level. To address this and to feed cleaner code for downstream semantic validation, D-Score adopts D-helix’s Recompiler. As an iterative recompilation tool, Recompiler automatically initializes undefined variables, injects required system libraries, and translates special pseudo-instructions (e.g., CONCAT) into function calls. For each candidate code $o_{i}$ from the baseline model, D-Score submits it to Recompiler and obtains a syntax score, $R_{syn}$, based on Recompiler’s feedback. Specifically, D-Score assign $R_{syn}(o_{i})$ to $syn_{pen}$ when the $o_{i}$ cannot be recompiled and passpass otherwise. Formally:
@@ IV-B1 Syntax Metric
Note that $syn_{pen}$ contributes directly to D-Score (See Equation˜3), when there is a syntax error, ensuring a clear penalty.
IV-B2 Semantics Metric
@@
Figure 5: The workflow of semantic check in D-Score. The left side shows how the function return value is checked, and the right side shows how the external function call is examined. For both scenarios, D-Score runs symbolic execution with symbolic inputs. After that, it constructs symbolic models, representing the function’s behaviors, by using If-Then-Else statements, which connect the inputs to the function’s return values/external calls.
@@
Once the candidate code $o_{i}$ passes the syntax check, D-Score runs symbolic execution to conduct the semantic check between the original binary and the intermediate representation (IR) code of $o_{i}$. To better describe, we split into two parts: return value and external function call.
@@
To confirm that the return values of the original binary and any candidate code match across different inputs, we mostly adopt D-helix’s approach, which employs symbolic execution and SMT solver. Specifically, as shown on the left side of Figure˜5, given the candidate code, D-Score first converts each function argument into a symbolic variable and performs symbolic execution to build a Symbolic-Model-Ret. This model captures the function’s semantics by linking symbolic inputs to arguments and symbolic outputs to return values, effectively encoding how each conditional transformation affects the final result. After that, D-Score runs the SMT solver on SMT-RET to compare the symbolic models between the original binary and the decompiled code.
@@
When verifying external function calls, however, directly applying D-helix to test LLM-generated candidate code (i.e., $o_{i}$) can yield many false negatives. This limitation stems from D-helix’s fundamental dependence on decompiled code to establish ground truth for verifying external function calls. Specifically, it models external function calls by approximating return values through the sum of the least significant bytes of those arguments. However, this external function call modeling process is only initialized when an external function call is detected within the decompiled code. Since instructions, including function calls, are often omitted in the LLM-generated code, D-helix may not properly initialize its function call modeling procedures, leading to possible false negatives. In fact, accurately restoring the prototypes for external function calls is an open problem [34], as this "ground truth" information is absent from the binary itself.
@@
Hence, to model the external function call in LLM-generated code in a better way, D-Score adopts a strategy that does not rely on unavailable prototype information. To do this, D-Score checks whether the number of external function calls being invoked, under varying input conditions, matches between the original binary and the LLM-generated code, as shown on the right side of Figure˜5. Specifically, to identify valid external calls in each decompiled function, we first extract and save the names of all invoked external functions to a file as our ground truth. After that, during the symbolic execution of the LLM-generated code, we build a symbolic model called Symbolic-Model-Call to model the external function call behavior. Similar to Symbolic-Model-Ret, this model is built through symbolic execution with symbolic inputs. However, instead of tracking return values, it counts how many ground-truth external function calls are invoked along each execution path. This count is then used as a behavioral signature of function-call activity. Finally, we run SMT solver on SMT-CALL to compare the function call-count symbolic models between the original binary and the decompiled code. By using this approach, we successfully avoid relying on potentially unreliable function prototype information.
By sequentially integrating the return-value check and the external-call check, we define our semantics metric as follows:
@@
$R_{sem}(o_{i},bin)=\left\{\begin{array}[]{l}ret_{pen},\text{ if }(check_{ret}(o_{i},bin)==false)\\ call_{pen},\text{ if }(check_{ret}(o_{i},bin)==true,\\ \quad\quad\quad\quad\text{and }check_{call}(o_{i},bin)==false)\\ pass,\text{ otherwise }\end{array}\right$. (8)
@@
, where $check_{ret}(o_{i},bin)$ is the boolean result of running SMT solver on SMT-RET, $check_{call}(o_{i},bin)$ is the result of running SMT solver on SMT-CALL, $ret_{pen}$ and $call_{pen}$ are the penalties applied when the return-value check or the external-call check fails, respectively.
IV-B3 Readability Metric